Let X be the thickness of washers a machine turns out. Assume that x has the density f(x)=2.5x if 1.9 < x < 2.1 and 0 otherwish then calculate the probability that washer will have thickness between 1.95mm and 2.05 mm.


1
Expert's answer
2021-05-07T09:44:47-0400
"P(1.9\leq X\leq 2.1)=\displaystyle\int_{1.95}^{2.05}2.5xdx"

"=[1.25x^2]\begin{matrix}
 2.05 \\
 1.95
\end{matrix}=1.25(2.05^2-1.95^2)"

"=0.5"